# The Edge-of-Reach Problem in Offline Model-Based Reinforcement Learning

### v1

@@ Abstract
Offline reinforcement learning aims to enable agents to be trained from pre-collected datasets, however, this comes with the added challenge of estimating the value of behavior not covered in the dataset. Model-based methods offer a solution by allowing agents to collect additional synthetic data via rollouts in a learned dynamics model. The prevailing theoretical understanding is that this can then be viewed as online reinforcement learning in an approximate dynamics model, and any remaining gap is therefore assumed to be due to the imperfect dynamics model. Surprisingly, however, we find that if the learned dynamics model is replaced by the true error-free dynamics, existing model-based methods completely fail. This reveals a major misconception. Our subsequent investigation finds that the general procedure used in model-based algorithms results in the existence of a set of edge-of-reach states which trigger pathological value overestimation and collapse in Bellman-based algorithms. We term this the edge-of-reach problem. Based on this, we fill some gaps in existing theory and also explain how prior model-based methods are inadvertently addressing the true underlying edge-of-reach problem. Finally, we propose Reach-Aware Value Learning (RAVL), a simple and robust method that directly addresses the edge-of-reach problem and achieves strong performance across both proprioceptive and pixel-based benchmarks.
Offline Reinforcement Learning, Model-Based Reinforcement Learning
@@ 1 Introduction
Figure 1: Boundaries of the states reachable at each timestep in the $k$-step rollouts from $\mathcal{D}_{\textrm{offline}}$ as used in offline model-based RL. In this paper we investigate a hitherto unrecognized pathology in offline model-based methods whereby edge-of-reach states - those only reachable at the final step of model rollouts (in red) - can trigger value overestimation and collapse of training.
@@
Standard online reinforcement learning (RL) requires the collection of large amounts of on-policy data throughout training. This can be both costly and unsafe, and hence represents a barrier against applying RL in domains such as healthcare (Tang & Wiens, 2021; Shiranthika et al., 2022) or robotics (Kumar et al., 2021; Chebotar et al., 2021; Ball et al., 2021), and also against scaling RL to more complex problems. Offline RL (Levine et al., 2020; Ernst et al., 2005) aims to remove this need for online data collection by enabling agents to be trained on pre-collected datasets. The use of large pre-existing datasets has driven significant recent advances in supervised learning (Brown et al., 2020), and one hope is that offline RL may provide a route for similar scaling in RL (Kumar et al., 2023).Code open-sourced at: github.com/anyasims/edge-of-reach.
@@
The central challenge in offline RL is dealing with value estimation for counterfactual actions not present in the dataset, termed the out-of-sample action problem (Kostrikov et al., 2022). [^1] Also referred to as the out-of-distribution action (Kumar et al., 2019) or action distribution shift problem (Kumar et al., 2020). A naïve approach results in inaccurate values at out-of-sample state-actions being propagated through the entire state-action space, leading to pathological training dynamics often exhibited as extreme value overestimation (Kumar et al., 2019). There have been many proposals to resolve this, with methods largely falling into one of two categories: model-free (Kostrikov et al., 2022; Kumar et al., 2020; Fujimoto et al., 2019; Kumar et al., 2019; An et al., 2021; Fujimoto & Gu, 2021) or model-based (Yu et al., 2020; Kidambi et al., 2020; Sun et al., 2023; Lu et al., 2022).
Model-free methods typically address this issue by applying a form of conservatism or constraint to avoid using out-of-sample actions in the Bellman update. In contrast, model-based methods, which are the focus of this paper, generally aim to resolve the issue by using the offline dataset to fit an approximate dynamics model (Sutton, 1991; Janner et al., 2019), and then allowing the agent to collect additional data in this learned dynamics model through fixed horizon rollouts from the original offline dataset (the general MBPO procedure (Janner et al., 2019), see Algorithm 1). The current understanding is that this converts the offline learning problem into online RL in an approximate dynamics model, thus solving the out-of-sample problem by allowing the agent to collect corrective feedback for any misestimated values. With this understanding, model-based methods typically attribute any remaining difficulties to errors in the learned dynamics model, and algorithms are subsequently motivated by avoiding exploitation of model errors using some notion of dynamics uncertainty.
This understanding of model-based approaches naturally implies that reducing dynamics model errors should improve performance. However, we begin with a surprising demonstration that existing offline model-based methods completely fail when trained with the true error-free dynamics model (see Table 1 below). This exposes a significant misconception in the current understanding of model-based offline RL and prompts us to reevaluate our understanding of how prior methods work. On investigation, we find that the use of short rollouts in these algorithms (as opposed to full horizon trajectories as collected in online RL) causes a critical and previously unnoticed problem. More specifically, it leads to the existence of states which, under any policy, can only be reached in the final rollout step (see red in Figure 1). These edge-of-reach states appear in the training distribution as nextstates, and are hence used in Bellman updates, but even with the ability to collect unlimited additional data, the agent is never able to observe the outcome of actions from them. The result is pathological overestimation of values, similar to that seen in model-free methods. We term this the edge-of-reach problem. Thus, contrary to common understanding, the out-of-sample problem central in model-free methods is not resolved by a model-based approach.
This edge-of-reach problem reveals strong connections between the two previously disjoint subfields of offline model-free and model-based RL. With this new understanding, we propose Reach-Aware Value Learning (RAVL), a robust scalable solution that takes an existing solution to the central out-of-sample problem in model-free literature, and directly applies it to the edge-of-reach problem in model-based methods (both proprioceptive and pixel-based). We first analyze RAVL on simple toy environments, and then show that it scales to the standard D4RL (Fu et al., 2020) and the V-D4RL (Lu et al., 2023) benchmarks. Furthermore, we anticipate that this unified understanding of model-based and model-free algorithms may pave the way for porting further strengths of existing model-free algorithms into model-based approaches.
Code is open-sourced at: github.com/anyasims/edge-of-reach
To summarize, our contributions are as follows:
  • We expose a significant misconception in the current understanding of offline model-based RL (Section 4.1). This leads to our hypothesis and definition of the edge-of-reach problem (Section 4.2). We provide evidence that this is the true underlying source of issues on the standard D4RL (Fu et al., 2020) benchmark (Section 4.4), before then constructing a simple environment which allows us to demonstrate and verify this hypothesis (Section 5).
  • Based on this, we propose Reach-Aware Value Learning (RAVL), a natural solution to the edge-of-reach problem (Section 6). We extensively evaluate RAVL across the D4RL and pixel-based V-D4RL (Lu et al., 2023) benchmarks (Section 7), including careful analysis into what RAVL is learning, demonstrating that it directly targets the edge-of-reach problem as intended (Section 7.1) and achieves state-of-the-art performance.
  • Finally, we give a more accurate explanation of the efficacy of prior methods, showing they unintentionally address the edge-of-reach problem (Section 7.4). [^2] Throughout the paper, please refer to Table 9 for the naming conventions used in experiments.
@@ 2 Background
Reinforcement Learning. We consider the standard RL framework (Sutton & Barto, 2018), in which the environment is formulated as a Markov Decision Process, $M=(\mathcal{S},\mathcal{A},T,R,\mu_{0},\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, $T(s^{\prime}|s,a)$ and $R(s,a)$ denote the transition and reward dynamics, $\mu_{0}$ the initial state distribution, and $\gamma\in(0,1)$ is the discount factor. The goal in reinforcement learning is to learn a policy $\pi(a|s)$ that maximizes the expected discounted return $\mathbb{E}_{\mu_{0},\pi,T}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$.
@@
Actor-Critic Algorithms. The broad class of algorithms we consider are actor-critic (Konda & Tsitsiklis, 1999) methods which jointly optimize a policy $\pi$ and state-action value function ($Q$-function). Actor-critic algorithms iterate between fitting the $Q$-function to the current policy (policy evaluation) and improving the policy relative to the $Q$-function (policy improvement). The policy evaluation step relies on approximate dynamic programming with Bellman updates, updating the $Q$-function using some dataset $\mathcal{D}$ as follows:

### v2

@@ Abstract
Offline reinforcement learning (RL) aims to train agents from pre-collected datasets. However, this comes with the added challenge of estimating the value of behaviors not covered in the dataset. Model-based methods offer a potential solution by training an approximate dynamics model, which then allows collection of additional synthetic data via rollouts in this model. The prevailing theory treats this approach as online RL in an approximate dynamics model, and any remaining performance gap is therefore understood as being due to dynamics model errors. In this paper, we analyze this assumption and investigate how popular algorithms perform as the learned dynamics model is improved. In contrast to both intuition and theory, if the learned dynamics model is replaced by the true error-free dynamics, existing model-based methods completely fail. This reveals a key oversight: The theoretical foundations assume sampling of full horizon rollouts in the learned dynamics model; however, in practice, the number of model-rollout steps is aggressively reduced to prevent accumulating errors. We show that this truncation of rollouts results in a set of edge-of-reach states at which we are effectively “bootstrapping from the void.” This triggers pathological value overestimation and complete performance collapse. We term this the edge-of-reach problem. Based on this new insight, we fill important gaps in existing theory, and reveal how prior model-based methods are primarily addressing the edge-of-reach problem, rather than model-inaccuracy as claimed. Finally, we propose Reach-Aware Value Learning (RAVL), a simple and robust method that directly addresses the edge-of-reach problem and hence - unlike existing methods - does not fail as the dynamics model is improved. Since world models will inevitably improve, we believe this is a key step towards future-proofing offline RL. [^1] Our code is open-sourced at: github.com/anyasims/edge-of-reach.
@@ 1 Introduction
Standard online reinforcement learning (RL) requires collecting large amounts of on-policy data. This can be both costly and unsafe, and hence represents a significant barrier against applying RL in domains such as healthcare [32, 28] or robotics [20, 4, 2], and also against scaling RL to more complex problems. Offline RL [23, 6] aims to remove this need for online data collection by enabling agents to be trained on pre-collected datasets. One hope [21] is that it may facilitate advances in RL similar to those driven by the use of large pre-existing datasets in supervised learning [3].
[…]
@@
The central challenge in offline RL is estimating the value of actions not present in the dataset, known as the out-of-sample action problem [17]. [^2] This is also referred to as the out-of-distribution action [18] or action distribution shift problem [19]. A naïve approach results in extreme value overestimation due to bootstrapping using inaccurate values at out-of-sample state-actions [18]. There have been many proposals to resolve this, with methods largely falling into one of two categories: model-free [17, 19, 10, 18, 1, 8] or model-based [36, 14, 29, 24].
Model-free methods typically address the out-of-sample action problem by applying a form of conservatism or constraint to avoid using out-of-sample actions in the Bellman update. In contrast, the solution proposed by model-based methods is to allow the collection of additional data at any previously out-of-sample actions. This is done by first training an approximate dynamics model on the offline dataset [30, 13], and then allowing the agent to collect additional synthetic data in this model via $k$-step rollouts (see Algorithm 1). The prevailing understanding is that this can be viewed as online RL in an approximate dynamics model, with the instruction being to then simply “run any RL algorithm on $\widehat{M}$ until convergence”[36], where $\widehat{M}$ is the learned dynamics model with some form of dynamics uncertainty penalty. Existing methods propose various forms of dynamics penalties (see Table 7), based on the assumption that the remaining performance gap compared to online RL is solely due to inaccuracies in the learned dynamics model.
This understanding naturally implies that improving the dynamics model should also improve performance. Surprisingly, we find that existing offline model-based methods completely fail if the learned dynamics model is replaced with the true, error-free dynamics model, while keeping everything else the same (see Figure 1). Under the true dynamics, the only difference to online RL is that in online RL, data is sampled as full-length episodes, while in offline model-based RL, data is instead sampled as $k$-step rollouts, starting from a state in the original offline dataset, with rollout length $k$ limited to avoid accumulating model errors. Failure under the true model therefore highlights that truncating rollouts has critical and previously-overlooked consequences.
We find that this rollout truncation leads to a set of states which, under any policy, can only be reached in the final rollout step (see red in Figure 2). The existence of these edge-of-reach states is problematic as it means Bellman updates (see Equation 1) use target values that are never themselves trained. This is illustrated in Figure 2, and described in detail in Section 3. This effective “bootstrapping from the void” triggers a catastrophic breakdown in $Q$-learning. Concisely, this issue can be viewed as all actions from edge-of-reach states remaining out-of-reach over training. Hence, contrary to common understanding, the out-of-sample action problem central to model-free methods is not fully resolved by a model-based approach. In fact, in Section 3.4 we provide detailed analysis suggesting that this is the predominant source of issues on the standard D4RL benchmark. In Section 6.5 we consequently reexamine how existing methods work and find they are indirectly and unintentionally addressing this issue (rather than model errors as claimed).
We have the following problem: Indirectly addressing the edge-of-reach problem means existing model-based methods have a fragile and unforeseen dependence on model quality and fail catastrophically as dynamics models improve. Since model improvements are inevitable, and since tuning of hyperparameters is particularly impractical in offline RL, this is a substantial practical barrier.
In light of this, we propose Reach-Aware Value Learning (RAVL), a simple and robust method that directly addresses the edge-of-reach problem. RAVL achieves strong performance on standard benchmarks, and moreover continues to perform even with error-free, uncertainty-free dynamics models (Section 6). Our model-based method bears close connections to model-free approaches, and hence in Section A.2 we present a unified perspective of these two previously disjoint subfields.
Algorithm 1: Pseudocode for the base procedure used in offline model-based methods.
In summary: (1) Train a dynamics model $\widehat{M}$ on $\mathcal{D}_{\textrm{offline}}$, then (2) Train an agent (often SAC) with $k$-step rollouts in the learned model starting from $s\in\mathcal{D}_{\textrm{offline}}$. Existing methods consider issues to be due to errors in $\widehat{M}$ and hence introduce dynamics uncertainty penalties (see Table 7). We find the predominant source of issues to be the edge-of-reach problem, and hence instead propose using value pessimism (RAVL) (see Section 5).
Algorithm 1 Base model-based algorithm (MBPO) + Additions in existing methods and RAVL (ours)
1:  Require: Offline dataset $\mathcal{D}_{\textrm{offline}}$
2:  Require: Dynamics model $\widehat{M}$ = ($\widehat{T}$, $\widehat{R}$) trained on $\mathcal{D}_{\textrm{offline}}$. Augment $\bm{\widehat{M}}$ with an uncertainty penalty *
3:  Specify: Rollout length $k\geq 1$, real data ratio $r\in[0,1]$
4:  Initialize: Replay buffer $\mathcal{D}_{\textrm{rollouts}}=\emptyset$, policy $\pi_{\theta}$, value function $Q_{\phi}$ (both from random)
5:  for $\textrm{epochs}=1,\dots$ do
6:     (Collect data) Starting from states in $\mathcal{D}_{\textrm{offline}}$, collect $k$-step rollouts in $\widehat{M}$ with $\pi_{\theta}$. Store data in $\mathcal{D}_{\textrm{rollouts}}$
7:     (Train agent) Train $\pi_{\theta}$ and $Q_{\phi}$ on $\mathcal{D}_{\textrm{rollouts}}\cup\mathcal{D}_{\textrm{offline}}$ (mixed with ratio $r$) Add $\mathbf{Q}$-value pessimism
8:  end for  *This uncertainty penalty collapses to zero with the true error-free dynamics ($\bm{\widehat{M}}=\bm{M}$) (see experiments Table 2).
@@ 2 Background
Reinforcement Learning. We consider the standard RL framework [31], in which the environment is formulated as a Markov Decision Process, $M=(\mathcal{S},\mathcal{A},T,R,\mu_{0},\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, $T(s^{\prime}|s,a)$ and $R(s,a)$ denote the transition and reward dynamics, $\mu_{0}$ the initial state distribution, and $\gamma\in(0,1)$ is the discount factor. The goal in reinforcement learning is to learn a policy $\pi(a|s)$ that maximizes the expected discounted return $\mathbb{E}_{\mu_{0},\pi,T}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$.
@@
Actor-Critic Algorithms. The broad class of algorithms we consider are actor-critic [15] methods which jointly optimize a policy $\pi$ and state-action value function ($Q$-function). Given a dataset $\mathcal{D}$ of (state, action, reward, nextstate) transitions, actor-critic algorithms iterate between two steps:
1. (Policy evaluation) Update $Q_{\phi}$ to fit the current policy using the following Bellman update: $\begin{split}\text{{\color[rgb]{0.8,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.0}\footnotesize{Key expression:}}}\qquad\underbrace{Q_{\phi}(s,a)}_{\textbf{\tiny{\color[rgb]{0.8,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.0}Update value at $\bm{(s,a)}$}}}\leftarrow\underbrace{r+\gamma Q_{\phi}(s^{\prime},\pi_{\theta}(s^{\prime}))}_{\textbf{\tiny{\color[rgb]{0.8,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.0}Using value at $\bm{(s^{\prime},\pi_{\theta}(s^{\prime}))}$}}}\qquad(s,a,r,s^{\prime})\sim\mathcal{D}\end{split}$ (1)
2. (Policy improvement) Update $\pi_{\theta}$ to increase the current $Q$-value predictions: $Q_{\phi}(s,\pi_{\theta}(s))$ [^3] This step contains an implicit $\max$ operation which we refer to in Section 3.2..
Offline RL and the Out-of-Sample Action Problem. In offline RL, training must rely on only a fixed dataset of transitions $\mathcal{D}_{\textrm{offline}}=\{(s,a,r,s^{\prime})^{(i)}\}_{i=1,\dots,N}$ collected by some policy $\pi^{\beta}$. The central problem in this offline setting is the out-of-sample [^4] We use the terms ‘out-of-sample’ and ‘out-of-distribution’ interchangeably, as is done in the literature. action problem: The values of $Q_{\phi}$ being updated are at any state-actions that appear as $(s,a)$ in $\mathcal{D}_{\textrm{offline}}$. However, the targets of these updates rely on values at $(s^{\prime},a^{\prime})$ where $a^{\prime}\sim\pi_{\theta}$ (see Equation 1). Since $a^{\prime}\sim\pi_{\theta}$ (on-policy) whereas $a\sim\pi^{\beta}$ (off-policy), $(s^{\prime},a^{\prime})$ may not appear in $\mathcal{D}_{\textrm{offline}}$, and hence may itself have never been updated. This means that updates can involve bootstrapping from ‘out-of-sample’ and hence arbitrarily misestimated values and thus lead to misestimation being propagated over the entire state-action space. The $\max$ operation (implicit in the policy improvement step) further acts to exploit any misestimation, converting misestimation into extreme pathological overestimation. As a result $Q$-values often tend to increase unboundedly over training, while performance collapses entirely [23, 18].
Model-Based Offline RL. Model-based methods [30] aim to solve the out-of-sample issue by allowing the agent to collect additional synthetic data in a learned dynamics model. They generally share the same base procedure as described in Algorithm 1. This involves first training an approximate dynamics model $\widehat{M}=(\mathcal{S},\mathcal{A},\widehat{T},\widehat{R},\mu_{0},\gamma)$ on $\mathcal{D}_{\textrm{offline}}$. Here, $\widehat{T}(s^{\prime}|s,a)$ and $\widehat{R}(s,a)$ denote the learned transition and reward functions, commonly realized as a deep ensemble [5, 22]. Following this an agent is trained using an online RL algorithm (typically SAC [11]), for which data is sampled as $k$-step trajectories (termed rollouts) under the current policy, starting from states in the offline dataset $\mathcal{D}_{\textrm{offline}}$. The base procedure of training a SAC agent with model rollouts does not work out of the box. Existing methods attribute this as due to dynamics model errors. Consequently, a broad class of methods propose augmenting the learned dynamics model $\widehat{M}$ with some form of dynamics uncertainty penalty, often based on variance over the ensemble dynamics model [36, 29, 14, 24, 2]. We present some explicit examples in Table 7. Crucially, all of these methods assume that, under the true error-free model, no intervention should be needed, and hence all the uncertainty penalties collapse to zero. In the following section, we show that this assumption leads to catastrophic failure.
Figure 2: The previously unnoticed edge-of-reach problem. Left illustrates the base procedure used in offline model-based RL, whereby synthetic data is sampled as $k$-step trajectories “rollouts” starting from a state in the original offline dataset. Edge-of-reach states are those that can be reached in $k$-steps, but which cannot (under any policy) be reached in less than $k$-steps. We depict the data collected with two rollouts, one ending in $s_{k}=D$, and the other with $s_{k}=C$. Right then shows this data arranged into a dataset of transitions as used in $Q$-updates. State $D$ is edge-of-reach and hence appears in the dataset as $s^{\prime}$ but never as $s$. Bellman updates therefore bootstrap from $D$, but never update the value at $D$ (see Equation 1). (For comparison consider state $C$: $C$ is also sampled at $s_{k}$, but unlike $D$ it is not edge-of-reach, and hence is also sampled at $s_{i<k}$ meaning it is updated and hence does not cause issues.)
3 The Edge-of-Reach Problem
In the following section, we formally introduce the edge-of-reach problem. We begin by showing the empirical failure of SOTA offline model-based methods on the true environment dynamics. Next, we present our edge-of-reach hypothesis for this, including intuition, empirical evidence on the main D4RL benchmark [7], and theoretical proof of its effect on offline model-based training.
3.1 Surprising Failure with the True Dynamics
As described in Sections 1 and 2, prior works view offline model-based RL as online RL in an approximate dynamics model. Based on this understanding they propose various forms of dynamics uncertainty penalties to address model errors (see Table 7). This approach is described simply as: “two steps: (a) learning a pessimistic MDP (P-MDP) using the offline dataset; (b) learning a near-optimal policy in this P-MDP”[14]. This shared base procedure is shown in Algorithm 1, where “P-MDP” refers to the learned dynamics model with a penalty added to address model errors.